OpenStudy (anonymous):

An electric field measured near a point charge of 3.00 × 10^-6 C has a strength of 1.50 × 10^5 N/C. How far is the point from the source charge? A.0.095 m B.0.30 m C.0.36 m D.0.42

OpenStudy (anonymous):

Actually can you help me on my question Michele_Laino

OpenStudy (michele_laino):

here w have to apply this formula: \[\large E = K\frac{Q}{{{d^2}}}\] so we have: \[\large d = \sqrt {\frac{{KQ}}{E}} \] where K=9*10^9, E is the electric field, and Q is the charge